The demand and supply equations for a product are:



Qd= 300 — 6P and Qs= -40 + 6P.



Determine the market equilibrium and draw graphs.


Suppose that the government decides to impose a flat tax of 10% on each unit sold. Show that the price that consumers pay would be the same if the government imposed a tax of Rs. 1.70 per unit sold. Draw graphs and explain.


Also calculate the total revenue earned by sellers before and after the tax, the tax revenue raised by the government, changes in consumer and producers surplus, and deadweight loss

1
Expert's answer
2021-12-16T16:19:13-0500

QS=QD

300 — 6P=-40 + 6P

12P=340

P=28.33

Q=300-6*28.33=130



after the tax is introduced, the supply function will be equal to

Qs=300-6(P-1.7)=289.8-6P



289.8-6P=-40 + 6P

12P=329.8

P=27.48

Q=124.9

The price has increased and the volume of sales has become less.